# A neural optimization framework for free-boundary diffeomorphic mapping problems and its applications

### v1

@@ 6 Spectral Beltrami Network
To overcome these limitations, we propose the Spectral Beltrami Network (SBN), a differentiable neural surrogate for $\mathcal{F}$ that enables gradient backpropagation through $\mu$ and $p$. The network operates on the unit disk $\mathbb{D}$, naturally following formulation 1. We represent the domain as a triangular mesh viewed as a directed graph $\mathcal{G}=(V,E,F)$ with nodes $V$, directed edges $E$, and faces $F$. SBN leverages three nested meshes $\mathcal{G}^{i}=(V^{i},E^{i},F^{i})$ $(i=1,2,3)$ of decreasing resolution. Given BC $\mu_{v}$ on all $v\in V^{1}$ and two points $p_{1},p_{2}\in V^{1}$, the network predicts a mapping $F$ approximating the numerical LSQC solution $f$ of $\{\mu_{T}\colon T\in F^{1},\mu_{T}=\frac{1}{3}\sum_{v\in T}\mu_{v}\}$ with fixed constraints $f(p_{i})=p_{i},i=1,2$. Enforcing selected two points to be fixed is based on two practical considerations. First, it constrains the output range, preventing training instabilities from unbounded target positions. Second, arbitrary target positions $q_{1},q_{2}$ can be recovered via post-composition with a similarity transformation $g$ satisfying $g(p_{j})=q_{j}$ $(j=1,2)$. By Prop.2, $\mathcal{F}(\mu,p_{1},p_{2},q_{1},q_{2})=g\circ\mathcal{F}(\mu,p_{1},p_{2},p_{1},p_{2})$, ensuring no loss of generality.
@@ 6.2 Message Passing Mechanism
We adapted the multiscale message passing from fortunato2022multiscale and improved the way of hierarchical edge construction. For downsampling graph $\mathcal{G}^{\text{down},i}=(V^{i}\cup V^{i+1},E^{i,i+1})$, each node $v_{k}\in V^{i}$ is connected to the three corners of its containing triangle in $F^{i+1}$ (identified by minimal $|\alpha_{k,l}|+|\alpha_{k,m}|+|\alpha_{k,n}|$ in barycentric coordinates $v_{k}=\alpha_{k,l}v_{l}+\alpha_{k,m}v_{m}+\alpha_{k,n}v_{n}$). Considering the free-boundary nature of LSQC, it is crucial to enhance boundary information exchange and thus we add edges from fine-mesh boundary nodes to any coarse-mesh vertex within a distance threshold. Upsampling graph $\mathcal{G}^{\text{up},i}=(V^{i}\cup V^{i+1},E^{i+1,i})$ reverses edge directions. See Fig.3. The general update Rules for node and edge embeddings are as follows: For graph $\mathcal{G}$ with edges $(v_{i},v_{j})$ where $v_{i}\in V^{k},v_{j}\in V^{l}$:
@@
Update Rules are applied sequentially to $\mathcal{G}^{k}$ (intra-mesh, $k=l$), $\mathcal{G}^{\text{down},k}$ (fine-to-coarse, $k<l$), and $\mathcal{G}^{\text{up},k}$ (coarse-to-fine, $k>l$). Here $f_{e},f_{v}$ are residual MLPs he2016deep with GraphNorm cai2021graphnorm; parameters $\theta_{e},\theta_{v}$ encode PDE constraints for robustness. This multiscale mechanism extends interaction ranges, allowing each node to reach the entire mesh through hierarchical edges and enhanced boundary broadcasting Fig.3 is the architecture of a message passing block.
@@ 6.3 Mesh Spectral Layer
Figure 2: Architecture of Message Passing Block.
Figure 3: Architecture of Mesh Spectral Layer
Inspired by Fourier Neural Operator (FNO)li2020fourier, we proposed the Mesh Spectral Layer (MSL) to capture global dependencies via the mesh Laplacian’s eigenbasis. For a triangular mesh $(V,E,F)$ with Laplacian $\Delta=\mathbf{M}^{-1}\mathbf{L}$, where
@@
where $\alpha_{ij}$ and $\beta_{ij}$ are the angles opposite to the edge $[i,j]$ and $\mathcal{N}(i)$ is the list of adjacent vertices and faces of the vertex $i$. Since $\Delta$ is symmetric positive semi-definite and encodes mesh geometry, we project latent features $X\in\mathbb{R}^{|V|\times d}$ onto its $k$ smallest eigenvectors $F\in\mathbb{R}^{k\times|V|}$, performing spectral-domain mixing:
@@
with trainable $W\in\mathbb{R}^{d\times\tilde{d}}$ and $R\in\mathbb{R}^{k\times\tilde{d}\times d}$. This offers efficient global communication with far fewer parameters than attention mechanisms vaswani2017attention. Fig.3 shows the architecture of this module.
@@ 6.4 Model architecture and training
A key principle in combining global operators with local message passing is that each node should carry structural information beyond individual messages. Fig.4 shows the complete SBN architecture integrating multiscale message passing and MSLs.
Figure 4: Architecture of Spectral Beltrami Network
SBN is trained by supervising the predicted mapping $\hat{u}=\mathcal{F}_{\theta}(\mu_{V^{1}},p_{1},p_{2})$ against ground-truth $u=\mathcal{F}_{\theta}(\mu_{F^{1}},p_{1},p_{2},p_{1},p_{2})$ where $\mu_{F^{1}}=\{\mu_{T}|\ \mu_{T}=\frac{1}{3}\sum_{v\in T}\mu_{v},T\in F^{1}\}$, via four loss components:
@@
    Boundary $L_{1}$ loss provides robust boundary fitting:
@@
In conclusion, the total training loss is a weighted sum of these loss terms. Once the training of SBN was finished, we obtained a neural network $\mathcal{F}_{\theta^{*}}$ and in the optimization problem, its parameters are frozen.
@@ 7 SBN-Opt: A novel optimization framework of the free-boundary diffeomorphism problem
setting $\mu_{v}=\mathcal{T}(\tilde{\mu}_{v},T_{\text{BC}})$ and $p_{i}=\mathcal{T}(\tilde{p}_{i},T_{\text{pin}})$ where $(\tilde{\mu}_{v},T_{\text{BC}},\tilde{p}_{i},T_{\text{pin}})$ are unconstrained optimization variables. In this way, we introduce the SBN guided optimization framework SBN-Opt for solving the problem 1
$\min_{(\tilde{\mu}_{v},T_{\text{BC}},\tilde{p}_{i},T_{\text{pin}},\phi,s,r)}E_{1}\Big(g\big(\mathcal{F}_{\theta^{*}}(\mu_{v},p_{1},p_{2}),\phi,s,r\big)\Big)+E_{2}(\mu),\quad g(x,\phi,s,r)=se^{i\phi}x+r.$
Direct optimization on BC enables explicit distortion control via $E_{2}(\mu)$. For instance, angle distortion is penalized by $E_{\text{BC}}=\frac{1}{|V^{1}|}\sum_{v\in V^{1}}|\mu_{v}|^{2}$, while smoothness (assuming piecewise-linear $\mu$) is enforced via $E_{\text{smooth}}=\frac{1}{|F^{1}|}\sum_{T\in F^{1}}|\nabla\mu(T)|^{2}$. Although SBN outputs only on $V^{1}$, any point $v\in\mathbb{D}$ maps via barycentric interpolation. By resolution-independence (Prop. 3), this remains faithful to the continuous solution. For each vertex in a high-resolution mesh $v_{p}\in V^{\text{orig}}$, we identify the triangle $I(p)=[v_{i},v_{j},v_{k}]\in F^{1}$ that minimizes $|\lambda_{i}|+|\lambda_{j}|+|\lambda_{k}|$ among barycentric representations $v_{p}=\lambda_{i}v_{i}+\lambda_{j}v_{j}+\lambda_{k}v_{k}$ (with $\lambda_{i}+\lambda_{j}+\lambda_{k}=1$) in all triangles, and this identifies the containing triangle where all coordinates are non-negative. We precompute a sparse interpolation matrix $R\in\mathbb{R}^{|V^{\text{orig}}|\times|V^{1}|}$ where
@@
Next, we take the tasks in Sec.9 as examples to illustrate how to apply the framework 7. In the density equalizing mapping problem, given a triangular mesh $\mathcal{T}=(V,F)$ of $\mathbb{D}$ with population $\{p(T):T\in F\}$, we seek a homeomorphism $f:\mathbb{D}\rightarrow\mathbb{R}^{2}$ making density $\rho(f(T))=\frac{p(T)}{\text{Area}(f(T))}$ uniform across faces. If without setting up a sea surrounding the disk, the region might arbitrary expand. However, setting up a sea might restrict the movement of the true boundary and make the area change near the true boundary damped. Instead, if needed, we can easily discourage arbitrary expansion of the region by designing a barrier function on the scaling parameter $s$ in the framework 7.
@@
where $\bar{\rho}(f)=\frac{\sum_{T}p(T)}{\sum_{T}\text{Area}(f(T))}$ is the target density. The total area is efficiently computed via Green’s theorem: $\sum_{T}\text{Area}(f(T))=\frac{1}{2}\sum_{[v_{i},v_{j}]\in\partial\mathbb{D}}\big(f(v_{i})_{1}f(v_{j})_{2}-f(v_{i})_{2}f(v_{j})_{1}\big)$.
In the work qiu2020inconsistent, Qiu et al. developed an efficient splitting method based on Alternating Direction Method with Multipliers(ADMM) to apply the numerical LSQC algorithm in the inconsistent surface registration problem. However, the algorithm relies on prescribed landmark point-to-point correspondence. One of advantages in our framework is that the pinned points itself are also optimization parameters therefore it can be applied to more general inconsistent surface registration problems. In practice, exact point to point correspondence is usually impractical and unreasonable. For one thing, manually labeling point to point correspondence is time-consuming. For another, measurement noise, meshing artifacts and smoothing alter vertex positions by millimeters thus treating those vertices as hard truths injects error. Therefore, it is more common to relax the correspondence to the relation between point sets and point sets. Given surfaces $S_{1},S_{2}$ with subsets $\{S_{i,j}\}$ and intensity functions $I_{i}:S_{i}\rightarrow\mathbb{R}$, we seek a diffeomorphism $f:S_{1}\rightarrow\mathbb{R}^{3}$ and optimal regions $\Omega_{i}^{*}\subset S_{i}$ minimizing:
@@
After conformally parameterizing $S_{i}$ to $\mathbb{D}$, the problem reduces to finding the optimal $f:\mathbb{D}\rightarrow\mathbb{R}^{2}$ under the maximality assumption $f(\Omega_{1})=\Omega_{2}=f(S_{1})\cap S_{2}$. We minimize $E_{1}=\lambda_{\text{I}}\,E_{\text{I}}+\lambda_{\text{pc}}\,E_{\text{pc}}$ while $E_{2}(\mu)$ provides distortion regularization. For the general optimization problem, we have the following algorithm.
@@
In practice, a disk conformal parameterization is not mandatory. To avoid unnecessary conformal distortion caused by the fixed boundary shape, sometimes we use the method in levy2002least to compute the free-boundary conformal parameterization of a surface. Again, due to Prop.3 and Prop.2, we can still compute the deformation by interpolation. The idea is that, let $\psi:S\rightarrow\mathbb{R}^{2}$ be a conformal parameterization of the surface $S$, $r_{0}=\max\limits_{v\in S}\|\psi(v)\|_{2}$ and $R$ is the sparse interpolation matrix of $f(S)$ with respect to the mesh $r_{0}\cdot\mathbb{D}=(r_{0}\cdot V^{1},F^{1})$, if $f:V^{1}\rightarrow\mathbb{R}^{2}$ is a deformation predicted by SBN, then deformation of $\psi(S)$ is $r_{0}\cdot Rf(V^{1})$.

### v2

@@ 5 Spectral Beltrami Network
With these advantages, LSQC should have been a perfect tool to derive a free-boundary quasiconformal map. However, there is a drawback in this method. Since $\mathcal{F}$ involves solving a huge sparse linear system, it is computationally expensive or impossible to backpropagate the gradient of loss function $\mathcal{L}$ w.r.t the mapping $\frac{\partial\mathcal{L}}{\partial f}$ to that w.r.t BC $\frac{\partial\mathcal{L}}{\partial\mu}$, destinations of pinned points $\frac{\partial\mathcal{L}}{\partial q_{i}}$, strictly following the chain rule. Moreover, the choice of the pinned points is a discrete manner and there lacks a clear mechanism to update it according to the loss values. The Alternating Direction Method of Multipliers (ADMM) partially mitigates this by updating $\mu$ based on the BCs reconstructed from the newly updated mapping $f$, but it does not address the lack of explicit gradients for pinned points $p_{i}$ and destinations $q_{i}$. Previous approaches, such as those by Qiu Qiu et al. (2019); Qiu and Lui (2020) circumvent this by imposing destinations of certain points in advance, which is often unrealistic in real-life scenarios.
To overcome these limitations, we propose the Spectral Beltrami Network (SBN), a differentiable neural surrogate for $\mathcal{F}$ that enables gradient backpropagation through $\mu$ and $p$. The network operates on the unit disk $\mathbb{D}$. This is a canonical choice: by the uniformization theorem, many surfaces of interest (after reducing to disk topology) admit a conformal parameterization to $\mathbb{D}$, and in practice there have been many efficient algorithms Choi and Lui (2015); DeLillo et al. (2004) for computing the disk or circular conformal parameterizations. We represent the domain as a triangular mesh viewed as a directed graph $\mathcal{G}=(V,E,F)$ with nodes $V$, directed edges $E$, and faces $F$. SBN leverages three nested meshes $\mathcal{G}^{i}=(V^{i},E^{i},F^{i})$ $(i=1,2,3)$ of decreasing resolution. Given BC $\mu_{v}$ on all $v\in V^{1}$ and two points $p_{1},p_{2}\in V^{1}$, the network predicts a mapping $F$ approximating the numerical LSQC solution $f$ of $\{\mu_{T}\colon T\in F^{1},\mu_{T}=\frac{1}{3}\sum_{v\in T}\mu_{v}\}$ with fixed constraints $f(p_{i})=p_{i},i=1,2$. Enforcing selected two points to be fixed is based on two practical considerations. First, it constrains the output range, preventing training instabilities from unbounded target positions. Second, arbitrary target positions $q_{1},q_{2}$ can be recovered via post-composition with a similarity transformation $g$ satisfying $g(p_{j})=q_{j}$ $(j=1,2)$. By Proposition 2, $\mathcal{F}(\mu,p_{1},p_{2},q_{1},q_{2})=g\circ\mathcal{F}(\mu,p_{1},p_{2},p_{1},p_{2})$, ensuring no loss of generality.
@@ 5.2 Message Passing Mechanism
We adopted the multiscale message passing from Fortunato et al. (2022) and improved the way of hierarchical edge construction. For downsampling graph $\mathcal{G}^{\text{down},i}=(V^{i}\cup V^{i+1},E^{i,i+1})$, each node $v_{k}\in V^{i}$ is connected to the three corners of its containing triangle in $F^{i+1}$ (identified by minimal $|\alpha_{k,l}|+|\alpha_{k,m}|+|\alpha_{k,n}|$ in barycentric coordinates $v_{k}=\alpha_{k,l}v_{l}+\alpha_{k,m}v_{m}+\alpha_{k,n}v_{n}$). Considering the free-boundary nature of LSQC, it is crucial to enhance boundary information exchange and thus we add edges from fine-mesh boundary nodes to any coarse-mesh vertex within a distance threshold. Upsampling graph $\mathcal{G}^{\text{up},i}=(V^{i}\cup V^{i+1},E^{i+1,i})$ reverses edge directions. See Figure 1. The general update Rules for node and edge embeddings are as follows: For graph $\mathcal{G}$ with edges $(v_{i},v_{j})$ where $v_{i}\in V^{k},v_{j}\in V^{l}$:
@@
Update Rules are applied sequentially to $\mathcal{G}^{k}$ (intra-mesh, $k=l$), $\mathcal{G}^{\text{down},k}$ (fine-to-coarse, $k<l$), and $\mathcal{G}^{\text{up},k}$ (coarse-to-fine, $k>l$). Here $f_{e},f_{v}$ are residual MLPs He et al. (2016) with GraphNorm Cai et al. (2021); parameters $\theta_{e},\theta_{v}$ encode PDE constraints for robustness. This multiscale mechanism extends interaction ranges, allowing each node to reach the entire mesh through hierarchical edges and enhanced boundary broadcasting. Figure 2(a) is the architecture of a message passing block.
@@ 5.3 Mesh Spectral Layer
Figure 2: Architecture of key modules in the Spectral Beltrami Network.
Inspired by Fourier Neural Operator (FNO)Li et al. (2020), we proposed the Mesh Spectral Layer (MSL) to capture global dependencies via the mesh Laplacian’s eigenbasis. For a triangular mesh $(V,E,F)$ with Laplacian $\Delta=\mathbf{M}^{-1}\mathbf{L}$, where
@@
where $\alpha_{ij}$ and $\beta_{ij}$ are the angles opposite to the edge $[i,j]$ and $\mathcal{N}(i)$ is the list of adjacent vertices and faces of the vertex $i$. Since $\Delta$ is symmetric positive semi-definite and encodes mesh geometry, we project latent features $X\in\mathbb{R}^{|V|\times d}$ onto its $k$ normalized eigenvectors $F\in\mathbb{R}^{k\times|V|}$ associated to the $k$ smallest eigenvalues, performing spectral-domain mixing:
@@
with trainable $W\in\mathbb{R}^{d\times\tilde{d}}$ and $R\in\mathbb{R}^{k\times\tilde{d}\times d}$. This offers efficient global communication with far fewer parameters than attention mechanisms Vaswani et al. (2017). Figure 2(b) shows the architecture of this module.
@@ 5.4 Model architecture and training
A key principle in combining global operators with local message passing is that each node should carry structural information beyond individual messages. Figure 3 shows the complete SBN architecture integrating multiscale message passing blocks and mesh spectral layers.
Figure 3: Architecture of Spectral Beltrami Network.
SBN is trained by supervising the predicted mapping $\hat{u}=\mathcal{F}_{\theta}(\mu_{V^{1}},p_{1},p_{2})$ against ground truth $u=\mathcal{F}_{\theta}(\mu_{F^{1}},p_{1},p_{2},p_{1},p_{2})$ where $\mu_{F^{1}}=\{\mu_{T}|\ \mu_{T}=\frac{1}{3}\sum_{v\in T}\mu_{v},T\in F^{1}\}$, via four loss components:
@@
    Boundary $\ell^{1}$ loss enhances boundary fitting:
@@
In conclusion, the total training loss is a weighted sum of these loss terms. Once the training of SBN was finished, we obtained a neural network $\mathcal{F}_{\theta^{*}}$ and during the optimization, its parameters are frozen.
@@ 6 SBN-Opt: A novel optimization framework of the free-boundary diffeomorphism problem
setting $\mu_{v}=\mathcal{T}(\tilde{\mu}_{v},T_{\text{BC}})$ and $p_{i}=\mathcal{T}(\tilde{p}_{i},T_{\text{pin}})$ where $(\tilde{\mu}_{v},T_{\text{BC}},\tilde{p}_{i},T_{\text{pin}})$ are unconstrained optimization variables. In this way, we introduce the SBN guided optimization framework SBN-Opt to solve the problem 1
$\min_{(\tilde{\mu}_{v},T_{\text{BC}},\tilde{p}_{i},T_{\text{pin}},\phi,s,r)}E_{1}\Big(g\big(\mathcal{F}_{\theta^{*}}(\mu_{v},p_{1},p_{2}),\phi,s,r\big)\Big)+E_{2}(\mu),\quad g(x,\phi,s,r)=se^{i\phi}x+r.$ (19)
A significant advantage of the SBN-Opt framework is that the pinned points $P=\{p_{1},p_{2}\}$ are treated as learnable parameters rather than unchanged constraints. While the optimization of these points occurs in a continuous coordinate space, the Spectral Beltrami Network requires input to coincide with the discrete vertices of the mesh. During optimization, updated coordinates $\tilde{p}$ are unlikely to align exactly with existing mesh nodes. To resolve this discrepancy while maintaining a differentiable pipeline, we employ a straight-through estimation strategy. In each forward pass, we first identify the mesh vertex $v_{near}\in V$ closest to the current continuous estimate $\tilde{p}$. The input to the network, $p_{in}$, is then constructed using a stop-gradient operator as follows:
$p_{in}=\text{stop\_gradient}(v_{near}-\tilde{p})+\tilde{p}$
This formulation ensures that 1. the valid vertex $v_{near}$ is passed as the input to the SBN; 2. during the backward pass, by the stop-gradient operator, the gradient of the loss function $\mathcal{L}$ with respect to the input flows directly to the continuous parameters, i.e. $\frac{\partial\mathcal{L}}{\partial\tilde{p}}=\frac{\partial\mathcal{L}}{\partial p_{in}}$.
Direct optimization on Beltrami coefficients enables explicit distortion control via $E_{2}(\mu)$. For instance, angle distortion is penalized by $E_{\text{BC}}=\frac{1}{|V^{1}|}\sum_{v\in V^{1}}|\mu_{v}|^{2}$, while smoothness (assuming piecewise-linear $\mu$) is enforced via $E_{\text{smooth}}=\frac{1}{|F^{1}|}\sum_{T\in F^{1}}|\nabla\mu(T)|^{2}$. Although SBN outputs only on $V^{1}$, any point $v\in\mathbb{D}$ maps via barycentric interpolation. By resolution-independence (Proposition 3), this remains faithful to the continuous solution. For each vertex in a high-resolution mesh $v_{p}\in V^{\text{orig}}$, we identify the triangle $I(p)=[v_{i},v_{j},v_{k}]\in F^{1}$ that minimizes $|\lambda_{i}|+|\lambda_{j}|+|\lambda_{k}|$ among barycentric representations $v_{p}=\lambda_{i}v_{i}+\lambda_{j}v_{j}+\lambda_{k}v_{k}$ (with $\lambda_{i}+\lambda_{j}+\lambda_{k}=1$) in all triangles, and this identifies the containing triangle where all coordinates are non-negative. We precompute a sparse interpolation matrix $R\in\mathbb{R}^{|V^{\text{orig}}|\times|V^{1}|}$ where
@@
Next, we take two image processing tasks in Section 8 as examples to illustrate how to apply the framework 19. In the Equiareal mapping problem, given a triangular mesh $\mathcal{T}=(V,F)$ of a surface $\mathcal{M}$ and its conformal parameterization $\phi:\mathcal{M}\to\mathbb{D}$ as well as population $\{p(T):T\in F\}$, we seek a homeomorphism $\psi:\mathbb{D}\to\mathbb{C}$ such that the mapping $f=\psi\circ\phi$ preserves the area of $\mathcal{T}$, or to say, makes density $\rho(f(T))=\frac{\text{Area}(T)}{\text{Area}(f(T))}$ uniform across faces. If without setting up a sea surrounding the diskChoi and Rycroft (2018), the region might arbitrary expand. However, setting up a sea might restrict the movement of the true boundary and make the area change near the true boundary damped. Instead, if needed, we can easily discourage arbitrary expansion of the region by designing a barrier function on the scaling parameter $s$ in the Framework 19. Then the task loss objective in the formulation 1 for this task is
@@
where $\bar{\rho}(f)=\frac{\sum_{T}\text{Area}(T)}{\sum_{T}\text{Area}(f(T))}$ is the target density.
In the work Qiu and Lui (2020), Qiu et al. developed an efficient splitting method based on Alternating Direction Method with Multipliers(ADMM) to apply the numerical LSQC algorithm in the inconsistent surface registration problem. However, the algorithm relies on prescribed landmark point-to-point correspondence. One of advantages in our framework is that the pinned points are also optimization parameters therefore it can be applied to more general inconsistent surface registration problems. In practice, exact point to point correspondence is usually impractical and unreasonable. For one thing, manually labeling point to point correspondence is time-consuming. For another, measurement noise, meshing artifacts and smoothing alter vertex positions by millimeters thus treating those vertices as hard truths introduces error. Therefore, it is more common to relax the correspondence to the relation between point sets and point sets. Given surfaces $S_{1},S_{2}$ with subsets $\{S_{i,j}\}$ and intensity functions $I_{i}:S_{i}\rightarrow\mathbb{R}$, we seek a diffeomorphism $f:S_{1}\rightarrow\mathbb{R}^{3}$ and optimal regions $\Omega_{i}^{*}\subset S_{i}$ minimizing:
@@
After conformally parameterizing $S_{i}$ to $\mathbb{D}$ by $\phi_{i}$, the problem reduces to finding the optimal $g=\phi_{2}\circ f\circ\phi_{1}^{-1}:\mathbb{D}\rightarrow\mathbb{R}^{2}$ under the maximality assumption $g\circ\phi_{1}(\Omega_{1})=\phi_{2}(\Omega_{2})=f(S_{1})\cap S_{2}$. We minimize $E_{1}=\lambda_{\text{I}}\,E_{\text{I}}+\lambda_{\text{pc}}\,E_{\text{pc}}$ while $E_{2}(\mu)$ provides distortion regularization. For the general optimization problem, the framework SBN-Opt is described in Algorithm 1.
@@
In practice, a disk conformal parameterization is not mandatory. To avoid unnecessary conformal distortion caused by the fixed boundary shape, sometimes we use the method in Lévy et al. (2002) to compute the free-boundary conformal parameterization of a surface. Again, due to Proposition 3 and Proposition 2, we can still compute the deformation by interpolation. The idea is that, let $\phi:\mathcal{M}\rightarrow\mathbb{R}^{2}$ be a conformal parameterization of the surface $\mathcal{M}$, $r_{0}=\max\limits_{v\in\mathcal{M}}\|\phi(v)\|_{2}$ and $R$ is the sparse interpolation matrix of $\phi(\mathcal{M})$ with respect to the mesh $r_{0}\cdot\mathbb{D}=(r_{0}\cdot V^{1},F^{1})$, if $f:V^{1}\rightarrow\mathbb{R}^{2}$ is a deformation predicted by SBN, then deformation of $\psi(S)$ is $r_{0}\cdot Rf(V^{1})$.